What are the implications of IS and LM curves? What are the factors on which the position and the slope of IS and LM curves depend?


1
Expert's answer
2021-05-13T10:49:51-0400

Implications of IS-LM curve-

An interest rate target can cause Y to change more than a money supply target when LM is set and IS moves left and right. An interest rate target holds Y steady while IS is set and LM swings left and right, but a money supply target (shifts in the LM curve) causes Y to swing wildly. 

This helps to understand why, in the 1970s and 1980s, many central banks abandoned money supply targeting in favor of interest rate targeting, during which time autonomous shocks to LM were common due to financial innovation, deregulation, and loophole mining.

One significant assumption is that if the IS curve becomes more unstable than the LM curve again, central banks will find it prudent to return to targeting monetary aggregates.

The IS-policy LM's power is severely restricted by its short-run expectation that the price level will remain constant.


IS is the interest rate and income equilibrium point in the asset market.

"Y=C(Y)+I(r)"

By doing total differentiation we get,

"\frac{dr}{dy}=\frac{(1-c)}{i}" , where c is equal to MPC and i is "\frac{di}{dr}"

As a result, the slope of IS is determined by MPC and i. When the government introduces a proportional tax at a rate of t, the slope is also determined by t.

Fiscal policy, on the other hand, determines its status.

Any change in government spending will move the IS, and any change in the factors that influence its slope will flatten or steepen the curve.

LM is the interest rate and income equilibrium point in the money market, i.e.

"\frac{M}{P}=L(Y,r)"

Similarly, we can calculate the slope of LM by differentiating the Ly and Lr, where Ly equals "\frac{dL}{dy }" and Lr equals "\frac{d}{dy}"

Its position is determined by monetary policy.